OpenStudy (anonymous):

A=1/2h(a+b), for b solve for the formula for the specefied variable

TheSmartOne (thesmartone):

\[A\frac{ 2 }{ h }=\frac{ 1 }{2 }h(a+b)\frac{ 2 }{h}\]

TheSmartOne (thesmartone):

\[\frac{ 2A }{ h }=a+b\]

TheSmartOne (thesmartone):

So now just subtract a on both sides and you have your answer...